# Second FRCSyn-onGoing: Winning Solutions and Post-Challenge Analysis to Improve Face Recognition with Synthetic Data

### v1

@@ V Second FRCSyn-onGoing: Results
This section provides an in-depth analysis of the results in terms of the different demographic groups and individual databases considered in the 2 FRCSyn-onGoing. Figure 12 shows the Detection Error Tradeoff (DET) curves of Sub-Tasks 1.1, 1.2, and 1.3, including the results achieved for the top-1 team in each demographic group. For completeness, the information and graphical representations for all the teams can be found on the challenge Codalab platform121212https://codalab.lisn.upsaclay.fr/competitions/16970#results.
@@
Figure 13: DET curves of Sub-Tasks 2.1 (left), 2.2 (Middle), and 2.3 (right), including the results achieved for the top-1 team (i.e., OPDAI, Idiap-SynthDistill, K-IBS-DS, respectively) in each evaluation database.
Figure 13 shows the DET curves of Sub-Tasks 2.1, 2.2, and 2.3, including the results achieved for the top-1 team in each database. Analyzing the FR model proposed by the OPDAI team for Sub-Task 2.1, the spread of the curves indicates variability in the system performance across different databases, with the results from AgeDB (94.54% Accuracy) outperforming others. Moreover, in Sub-Task 2.2 the Idiap-SynthDistill FR model significantly improves the performance of Sub-Task 2.1 for AgeDB and CFP-FP databases (i.e., 96.72% and 96.14% Accuracy, respectively). Finally, for Sub-Task 2.3, the curves from the K-IBS-DS FR model are closely aligned for the AgeDB (96.89% Accuracy), BUPT (96.88% Accuracy), and CFP-FP (97.51% Accuracy), showing consistent and reliable performance across these databases. However, the curve from the ROF database remains the worst in each sub-task (i.e., 87.22%, 87.31%, and 90.39% Accuracy for Sub-Tasks 2.1, 2.2, and 2.3, respectively), reflecting that it is the most difficult database to emulate with synthetic data.
@@
Analyzing the contributions of all eleven top teams, we can observe the prevalence of well-established methodologies. Notably, most teams used DCFace [28] either independently or in conjunction with other synthetic databases such as GANDiffFace [21], DigiFace-1M [31], or IDiff-Face [29]. Furthermore, several teams, including CTAI and CBSR-Samsung, adopted interesting approaches involving synthetic data cleaning and selection. It is interesting to highlight the ID R&D and Idiap-SynthDistill teams as they considered novel methods to generate synthetic data. Specifically, the ID R&D team used an HDT [54] to generate synthetic facial images along with identity and style embeddings, wich were used by a StyleNAT [56] model to generate more variability in the synthetic data. Another example is the Idiap-SynthDistill team, which proposed an end-to-end method that dynamically generated facial images through StyleGAN2 [62] and trained a FR model through model distillation. Regarding the backbone architecture, all teams opted for either ResNet [36] or IResNet [48], mainly for their widespread adoption in state-of-the-art FR methodologies. The selection of the loss functions was also similar among the teams, with AdaFace [7] and ArcFace [6] the prevalent choices. Nevertheless, there were exceptions such as for the ID R&D team that used the recent UniFace [58], or the UNICA team that considered CosFace [61].
TABLE VII: Description of the best results achieved in the 1 and 2 FRCSyn-onGoing. Sub-Tasks 1.2 and 2.2 of the 2 FRCSyn-onGoing are not included in the table as they are novel sub-tasks only available in the 2 edition.
@@ 2 FRCSyn-onGoing
Sub-Task 1.3: Bias Mitigation
1 FRCSyn-onGoing
2 FRCSyn-onGoing
Team
TO [%]
GAP [%]
Team
TO [%]
GAP [%]
CBSR
95.25
-2.10
ADMIS
96.50
-1.33
@@
Sub-Task 2.3: Overall Improvement
1 FRCSyn Challenge
2 FRCSyn-onGoing
Team
AVG [%]
GAP [%]
Team
AVG [%]
GAP [%]
BOVIFOCR
94.95
-3.69
K-IBS-DS
95.42
-2.15
Finally, we compare the results achieved in the 2 FRCSyn-onGoing with the results of the 1 edition [34]. Table VII shows the best results achieved in the 1 and 2 editions of the challenge, including also the GAP values. It is important to remark that Sub-Tasks 1.2 and 2.2 of the 2 FRCSyn-onGoing are not included in the analysis as they are novel sub-tasks only available in the 2 edition. Notably, two observations can be made: i) the main metric for ranking teams (i.e., TO and AVG) shows improvements across all cases in this 2 edition for both Task 1 (e.g., 96.73% vs. 92.25% TO in Sub-Task 1.1 and 96.50% vs. 95.25% TO in Sub-Task 1.3) and Task 2 (e.g., 91.93% vs. 90.50% AVG in Sub-Task 2.1 and 95.42% vs. 94.95% AVG in Sub-Task 2.3), and ii) in terms of the GAP value, the FR models of this 2 edition follow a similar trend compared to the 1st edition, achieving in most sub-tasks negative GAP values, remarking the benefits for training using synthetic data. In particular, for Sub-Task 1.1, a much higher negative GAP value is observed in this 2 edition (i.e., -5.31% vs. -0.74%). This result, together with the higher TO value, seems to be due to the generation of better synthetic data by the ID R&D team, with the proposal of novel generative methods, as indicated before. In addition, several conclusions can be drawn. First, the improvement in the main metric can be associated with the freedom to select the methodology to generate the synthetic data to train the FR models, as well as the application of data cleaning and selection techniques. Finally, the increasing GAP value can be associated also with the enhancement of the FR models, due to the proposal of different architectures and loss functions, as indicated before.
@@ Acknowledgements
This study has received funding from the European Union’s Horizon 2020 TReSPAsS-ETN (No 860813) and is supported by INTER-ACTION (PID2021-126521OB-I00 MICINN/FEDER), Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2), R&D Agreement DGGC/UAM/FUAM for Biometrics and Cybersecurity, and PowerAI+ (SI4/PJI/2024-00062, funded by Comunidad de Madrid through the grant agreement for the promotion of research and technology transfer at UAM).
It is also supported by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.
K-IBS-DS was supported by the Institute for Basic Science, Republic of Korea (IBS-R029-C2). UNICA-IGD-LSI was supported by the ARIS program P2-0250B.

### v2

@@ V-C Demographic Groups and Evaluation Databases
This section provides an in-depth analysis of the results in terms of the different demographic groups and individual databases considered in the 2 FRCSyn-onGoing. Figure 12 (left), shows the Detection Error Tradeoff (DET) curves of Sub-Tasks 1.1, 1.2, and 1.3, including the results achieved for the top-1 team in each demographic group. For completeness, the information and graphical representations for all the teams can be found on the challenge Codalab platform121212https://codalab.lisn.upsaclay.fr/competitions/16970#results.
@@
Figure 12 (right) shows the DET curves of Sub-Tasks 2.1, 2.2, and 2.3, including the results achieved for the top-1 team in each database. Analyzing the FR model proposed by the OPDAI team for Sub-Task 2.1, the spread of the curves indicates variability in the system performance across different databases, with the results from AgeDB (94.54% Accuracy) outperforming others. Moreover, in Sub-Task 2.2 the Idiap-SynthDistill FR model significantly improves the performance of Sub-Task 2.1 for AgeDB and CFP-FP databases (i.e., 96.72% and 96.14% Accuracy, respectively). Finally, for Sub-Task 2.3, the curves from the K-IBS-DS FR model are closely aligned for the AgeDB (96.89% Accuracy), BUPT (96.88% Accuracy), and CFP-FP (97.51% Accuracy), showing consistent and reliable performance across these databases. However, the curve from the ROF database remains the worst in each sub-task (i.e., 87.22%, 87.31%, and 90.39% Accuracy for Sub-Tasks 2.1, 2.2, and 2.3, respectively), reflecting that it is the most difficult database to emulate with synthetic data.
@@ V-D Post-Challenge Analysis and Comparison with 1 Edition
Analyzing the contributions of all eleven top teams, we can observe the prevalence of well-established methodologies. Notably, most teams used DCFace [22] either independently or in conjunction with other synthetic databases such as GANDiffFace [21], DigiFace-1M [29], or IDiff-Face [23]. DCFace is a Dual Condition Face Generator based on a diffusion model, designed to create facial images of the same subject in various styles while maintaining identity consistency. Its key component is the Patch-wise Style Extractor, which extracts style features from an image while minimizing identity information. This forces the model to rely on a separate input for identity data. Unlike previous approaches like StyleGAN [69], DCFace retains essential spatial details, such as pose, ensuring greater variability between subjects of the same identity. This results in images with a similar style that enhance the performance of facial recognition models in identifying subjects. Furthermore, several teams, including CBSR-Samsung, INESC-IGD, and CTAI adopted interesting approaches involving synthetic data cleaning and selection. These approaches include: i) de-overlapping the data from DCFace, as it is trained with CASIA-WebFace and some data could be very similar, deteriorating the training, ii) balancing the data with respect to the demographic information, and iii) removing the images that are far from the class center using clustering techniques such as DBSCAN. It is interesting to highlight the ID R&D and Idiap-SynthDistill teams as they considered novel methods to generate synthetic data. Specifically, the ID R&D team used an HDT [61] to generate synthetic facial images along with identity and style embeddings, which were used by a StyleNAT [63] model to generate more variability in the synthetic data. Another example is the Idiap-SynthDistill team, which proposed an end-to-end method that dynamically generated facial images through StyleGAN2 [70] and trained a FR model through model distillation. Regarding the backbone architecture, all teams opted for either ResNet [34] or IResNet [57], mainly for their widespread adoption in state-of-the-art FR methodologies. Both architectures use residual connections to improve the training, but while in ResNet, the skip connections bypass one or more layers to address the vanishing gradient problem, IResNet optimizes the information flow through the network, allowing for the training of extremely deep architectures without increasing model complexity. Finally, the selection of the loss functions was also similar among the teams, with AdaFace [7] and ArcFace [6] the prevalent choices. Both losses use angular margin-based loss functions to improve facial feature discrimination, but while in ArcFace the margin is fixed for all samples, AdaFace adapts the margin dynamically based on the quality of each image. Nevertheless, there were exceptions such as for the ID R&D team that used the recent UniFace [65], or the UNICA team that considered CosFace [68].
@@ 2 FRCSyn-onGoing
Sub-Task 1.3: Bias Mitigation 1 FRCSyn-onGoing 2 FRCSyn-onGoing Team TO [%] GAP [%] Team TO [%] GAP [%] CBSR 95.25 -2.10 ADMIS 96.50 -1.33 |
@@
Sub-Task 2.3: Overall Improvement 1 FRCSyn Challenge 2 FRCSyn-onGoing Team AVG [%] GAP [%] Team AVG [%] GAP [%] BOVIFOCR 94.95 -3.69 K-IBS-DS 95.42 -2.15 |
Aditionally, we compare the results achieved in the 2 FRCSyn-onGoing with the results of the 1 edition [32]. Table X shows the best results achieved in the 1 and 2 editions of the challenge, including also the GAP values. It is important to remark that Sub-Tasks 1.2 and 2.2 of the 2 FRCSyn-onGoing are not included in the analysis as they are novel sub-tasks only available in the 2 edition. Notably, two observations can be made: i) the main metric for ranking teams (i.e., TO and AVG) shows improvements across all cases in this 2 edition for both Task 1 (e.g., 96.73% vs. 92.25% TO in Sub-Task 1.1 and 96.50% vs. 95.25% TO in Sub-Task 1.3) and Task 2 (e.g., 91.93% vs. 90.50% AVG in Sub-Task 2.1 and 95.42% vs. 94.95% AVG in Sub-Task 2.3), and ii) in terms of the GAP value, the FR models of this 2 edition follow a similar trend compared to the 1st edition, achieving in most sub-tasks negative GAP values, remarking the benefits for training using synthetic data. In particular, for Sub-Task 1.1, a much higher negative GAP value is observed in this 2 edition (i.e., -5.31% vs. -0.74%). This result, together with the higher TO value, seems to be due to the generation of better synthetic data by the ID R&D team, with the proposal of novel generative methods, as indicated before. In addition, several conclusions can be drawn. First, the improvement in the main metric can be associated with the freedom to select the methodology to generate the synthetic data to train the FR models, as well as the application of data cleaning and selection techniques. We observe that the increasing GAP value can be associated also with the enhancement of the FR models, due to the proposal of different architectures and loss functions, as indicated before.
Finally, we observe that when it comes to the quality of the generated synthetic data, higher quality does not imply better performance in recognition tasks. Most teams use DCFace as the main dataset for their training, even though it generates images with lower resolution and less detail from a qualitative perspective. Nevertheless, analyzing the results achieved using this database, we conclude that FR models might not require highly detailed images to learn to match identities, at least for the FR databases considered in this challenge. This suggests that instead of focusing on realism, synthetic data for FR should generate diverse images that help models better learn class variability while reducing noise.
@@ Acknowledgements
This study is supported by INTER-ACTION (PID2021-126521OB-I00 MICINN/FEDER), Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2), R&D Agreement DGGC/UAM/FUAM for Biometrics and Cybersecurity, and PowerAI+ (SI4/PJI/2024-00062, funded by Comunidad de Madrid through the grant agreement for the promotion of research and technology transfer at UAM). It is also supported by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science, and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. K-IBS-DS was supported by the Institute for Basic Science, Republic of Korea (IBS-R029-C2). UNICA-IGD-LSI was supported by the ARIS program P2-0250B.